Given that Kw for water is 2.4× 10–14 M2 at 37 °C, compute the pH of a neutral aqueous solution at 37 °C, which is the normal human body temperature.

PH?
Answer #1

Kw = [H3O+][OH-]

2.4*10^-14 = x*x

x = [H3O+] = [OH-] = 1.55*10^-7 M

pH = -log(H3O+)

     = -log(1.55*10^-7)

   = 6.81